Find the slopes of the tangents to the curves obtained by slicing the surface

x^{2}e^{y}-yze^{x}=0 with planes x=1 and y=1 at the point (1, 1), using implicit

differentiation.


1
Expert's answer
2021-11-30T09:57:17-0500

"x^2e^y-yze^x=0"


"x^2e^y=yze^x"


Differentiate Z with respect to x to get the first slope tangent"\frac{dz}{dx}"


"\frac{dz}{dx}=\frac{2xe^{y-x}-zy}{y}"


Replace (1, 1,0) to get the slope since at z plane z=0

"\frac{dz}{dx}=2"


The first slope is "2"



The second slope

Differentiate z with respect to y


"\frac{dz}{dy}\\"


"\frac{dz}{dy}=\frac{x^2e^{y-x}-z}{y}"


Replace (1, 1,0) in the equation


"\frac{dz}{dy}=1"


The second slope is "1"